# Percolation as a confinement order parameter in $\mathbb{Z}_{2}$ lattice gauge theories

### 16 Jan 2024 v1

@@ Abstract
Lattice gauge theories (LGTs) were introduced in 1974 by Wilson to study quark confinement. These models have been shown to exhibit (de-)confined phases, yet it remains challenging to define experimentally accessible order parameters. Here we propose percolation-inspired order parameters (POPs) to probe confinement of dynamical matter in $\mathbb{Z}_{2}$ LGTs using electric field basis snapshots accessible to quantum simulators. We apply the POPs to study a classical $\mathbb{Z}_{2}$ LGT and find a confining phase up to temperature $T=\infty$ in 2D (critical $T_{c}$, i.e. finite-$T$ phase transition, in 3D) for any non-zero density of $\mathbb{Z}_{2}$ charges. Further, using quantum Monte Carlo we demonstrate that the POPs reproduce the square lattice Fradkin-Shenker phase diagram at $T=0$ and explore the phase diagram at $T>0$. The correlation length exponent coincides with the one of the 3D Ising universality class and we determine the POP critical exponent characterizing percolation. Our proposed POPs provide a geometric perspective of confinement and are directly accessible to snapshots obtained in quantum simulators, making them suitable as a probe for quantum spin liquids.
Introduction.–Lattice gauge theories (LGTs) have been widely studied in the fields of high-energy Wilson (1974), condensed matter Wegner (1971); Fradkin and Susskind (1978a); Kogut (1979) and biophysics Lammert et al. (1993). In 1979, Fradkin and Shenker proved in their groundbreaking work Fradkin and Shenker (1979) the existence of two phases in their model, where $\mathbb{Z}_{2}$ charged particles are confined or deconfined, respectively. Ever since, researchers have found intimate connections of $\mathbb{Z}_{2}$ LGTs to other physical effects, including topological order Wen (2007); Verresen et al. (2022), quantum spin liquids Read and Sachdev (1991); Sachdev (2019) and even quantum information Kitaev (2003). In the light of quantum simulation, LGTs with finite-dimensional local Hilbert spaces, e.g. $\mathbb{Z}_{2}$ LGTs and quantum link models Wiese (2013), gain particular attention because of their experimental feasibility Halimeh et al. (2023).
@@
Site percolation theory was first introduced in works by Flory and Stockmayer in the 1940s Flory (1941); Stockmayer (1944), where they studied the gel-point and the cross-linking of polymers. Nearly two decades later, bond percolation theory was introduced by Broadbent and Hammersley Broadbent and Hammersley (1957). They studied the random flow of a fluid through a medium and introduced the so-called Bernoulli percolation. A prime example of the use of percolation theory in physics is the random cluster (RC) model, which was introduced by Fortuin and Kasteleyn Kasteleyn and Fortuin (1969); Fortuin and Kasteleyn (1972); Fortuin (1972a, b). The RC model was successfully used to study many physical systems, e.g. the Potts model Grimmett (2006); Baxter (1982). More recently, percolation theory was applied in the context of quantum monopole motion in spin ice Stern et al. (2021).
In this Letter we show that percolation can also serve as a non-local order parameter to probe confinement. In the confined phase, analogous to quarks in a confining potential, two charges are being held together by $\mathbb{Z}_{2}$ electric strings and form a $\mathbb{Z}_{2}$-neutral meson Homeier et al. (2023). In the deconfined phase, in our picture matter particles move essentially independently through the system and the $\mathbb{Z}_{2}$ electric strings form a global cluster spanning over the entire lattice, see Fig. 1. Charges attached to this cluster are no longer constrained to $\mathbb{Z}_{2}$ neutral hadrons. We study a classical and a quantum model using classical (quantum) Monte Carlo (MC/QMC). We show that the phase boundary in the celebrated Fradkin-Shenker model (extended toric code) Fradkin and Shenker (1979) can be reproduced by percolation-inspired order parameters (POPs), paving the way for the application of POPs in related models.
@@
Classical $\mathbb{Z}_{2}$ LGT.– The complex nature of percolation in $\mathbb{Z}_{2}$ LGTs emerges from the local $\mathbb{Z}_{2}$ symmetry generator
@@
To probe this transition, we define two POPs. Percolation probability is the probability that a cluster percolates, i.e. winds around the system and connects to itself, see Sup Sec. I.1. In the literature, this concept is sometimes also referred to as the wrapping probability Feng et al. (2008). Percolation strength extends the percolation probability to include information about the size of the percolating cluster. When the system is not percolating, the percolation strength is zero. When the system is percolating, the percolation strength is defined as the number of strings in the largest string cluster divided by the total number of links Homeier et al. (2023). We probe the transition on a periodic square lattice with system sizes $L^{2}=20^{2},30^{2},40^{2},50^{2}$. In the following we distinguish the cases for (i) zero matter density $d=0$ and (ii) non-zero matter density $d=N/L^{2}>0$.
@@
Quantum $\mathbb{Z}_{2}$ LGT: extended toric code.–We start with a generalization of Hamiltonian (Percolation as a confinement order parameter in $\mathbb{Z}_{2}$ lattice gauge theories) by adding magnetic fluctuations $\propto J$ and dynamical matter $\propto\lambda$:
@@
At $h=\lambda=0.2$ (deconfined, topological ground state), we find a percolating phase at low but finite temperatures and a non-percolating phase at higher temperatures. We trace this back to the bulk gap, where due to exponential suppression of thermal matter excitations the topological ground state stays robust against a non-zero temperature in finite-size systems, see Sup Sec. II.2. At temperatures well above the bulk gap, we observe a clear decrease in the percolation strength with higher system sizes. We conjecture that the percolation probability approaches zero in the TDL for any $T>0$. We were unable to provide a good estimate of the Fredenhagen-Marcu order parameter for finite temperatures due to high noise in the data.
@@
Discussion and outlook.– We have introduced new POPs to probe the confinement of matter excitations in $\mathbb{Z}_{2}$ LGTs. We have applied it to study a classical model where we demonstrated the substantial influence of matter and the lattice geometry on the existence of a percolating, i.e. deconfined, phase. Further, using snapshots generated from continuous-time QMC, we demonstrated that the phase boundaries of the square lattice Fradkin-Shenker model, one of the textbook examples featuring confinement and topological order, can be reproduced by our POPs. The correlation length critical exponent falls into the 3D Ising universality class. We also demonstrated the potential of the POPs at finite temperature. Another task, left for future research, will be to check for a first-order percolation transition across the self-duality line at the tip of the deconfined phase.
While we only discuss $\mathbb{Z}_{2}$ LGTs on the square and the cubic lattice, we also envision future applications of the POPs in other LGTs, where different interactions, lattice geometries or higher symmetries can be realized. Percolation was already used in the context of quantum error correcting codes Botzung et al. (2023). Since percolation can be directly measured from snapshots in the electric field basis, the order parameter is readily accessible to state-of-the-art quantum simulators Zohar et al. (2017); Barbiero et al. (2019); Schweizer et al. (2019); Homeier et al. (2021, 2023); Mildenberger et al. (2022); Halimeh et al. (2023).
We thank M. Aidelsburger, G. Dünnweber, M. Kebrič, F. Palm, G. Semeghini, T. Wang and H. Zhao for fruitful discussions. This research was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement no 948141) — ERC Starting Grant SimUcQuam, and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868 and via Research Unit FOR 2414 under project number 277974659. L.H. acknowledges support from the Studienstiftung des deutschen Volkes.
@@ I.2 Grand Canonical Monte Carlo Sampling
$\mathcal{H}_{\mathrm{gc}}=-h\sum_{l}\hat{\tau}^{x}_{l}-\mu\,\sum_{\bm{j}}\,\frac{1}{2}\,\biggl{(}1-\prod_{l\in+_{\bm{j}}}\hat{\tau}^{x}_{l}\biggr{)}.$ (S5)
@@ II.2 Extended toric code QMC
$h$-sweeps.– We set $T=1/L$, $\lambda=0,0.1,0.2,0.3$ and $L^{2}=20^{2},25^{2},30^{2}$. To extract the critical electric field $h_{c}$ and the critical exponents $\nu$ (correlation length) and $\beta$ (percolation strength $P$) we first perform a crossing point analysis of the percolation strength Binder cumulant
@@
$\lambda$-sweep.– We set $T=1/L$, $h=0.2$ and $L^{2}=10^{2},20^{2},30^{2}$. We find Binder cumulant crossing points around $\lambda\approx 0.4$ that shift toward lower $\lambda$ with increasing system size, see Fig. S5c. The gap consists of the bulk gap $\Delta_{\mathrm{B}}$ and the finite-size gap $\Delta_{L}$. Because of the self-duality, we know that the true phase transition where the gap closes is at $\lambda_{c}\sim 0.33$. However, in finite-size systems (i) the gap never completely closes and (ii) the finite-size ground state can appear topological because of the finite-size gap. We conjecture that the true critical value of $\lambda_{c}\sim 0.33$ is reached in the TDL and the strong finite-size effects are caused by a large finite-size gap. In Fig. S5b, we show the Fredenhagen-Marcu order parameter
@@
Figure S6: Ising mapping of the pure gauge model. We illustrate the mapping of the pure gauge model Hamiltonian (S11) to the TFIM on the dual lattice. In panel (a) we show two configurations of electric strings and the corresponding spin configuration on the dual lattice. The electric strings form closed loops. The electric field on a link $l$ depends on the dual lattice link between $\bm{i}^{\prime}$ and $\bm{j}^{\prime}$ that intersects $l$. By flipping a spin on a dual lattice site $\bm{j}^{\prime}$, all electric fields in the surrounding plaquette $\square_{\bm{j}^{\prime}}$ are flipped. A percolating string corresponds to a domain wall of spins since closed string loops necessarily have spins of the same orientation along their perimeter. Due to Gauss’s law, the star term $\propto\mu$ is constant, see panel panel (b). Adapted from Wegner (2014).
$T$-sweep.– We set $h=\lambda=0.2$ and $L^{2}=10^{2},20^{2},30^{2}$. We find a crossing-point in the percolation strength Binder cumulant at $T>0$ and a percolating phase extending into $T>0$, see Fig. S5d. However, the topological order is known to break down for $T>0$. We trace this behavior back to the bulk gap $\Delta_{\mathrm{B}}$. For $T<\Delta_{\mathrm{B}}$ the density of thermal excitations is non-zero in the TDL, however, it is exponentially suppressed and thus not visible in finite-size systems. After all, this is the reason why we can gain insights into the ground state using finite-T QMC schemes. While the true ground state for $T>0$ is not topological, it appears like the topological ground state at $T=0$ in finite-size systems. This reasoning is supported by the estimated order-of-magnitude of the bulk gap $\Delta_{\mathrm{B}}\simeq J$ at $h=\lambda=0.2$ which matches the temperature scale below which we observe the percolating (deconfined) ground state features. We were unable to provide a good estimate of the Fredenhagen-Marcu order parameter for finite temperatures due to high noise in the data.
@@ III Mapping the pure gauge theory to the 2D transverse-field Ising model
Since we have no matter, the $\mathbb{Z}_{2}$ electric strings can only form closed loops. If we define spins $\hat{s}^{z}\pm 1$ on the dual lattice sites, a string loop on an elementary plaquette can be thought of as one spin with $\hat{s}^{z}=+1$ in a background of spins with $\hat{s}^{z}=-1$. The operator $\hat{s}^{x}$ flips the spin $\hat{s}^{z}$ in the dual lattice and thus the electric fields on the surrounding plaquette in the LGT: it creates/destroys string loops. This intuition can be formalized with the definition of the dual-lattice site variables:

### 08 Mar 2024 v2

@@ Abstract
Lattice gauge theories (LGTs) were introduced in 1974 by Wilson to study quark confinement. These models have been shown to exhibit (de)confined phases, yet it remains challenging to define experimentally accessible order parameters. Here we propose percolation-inspired order parameters (POPs) to probe confinement of dynamical matter in $\mathbb{Z}_{2}$ LGTs using electric field basis snapshots accessible to quantum simulators. We apply the POPs to study a classical $\mathbb{Z}_{2}$ LGT and find a confining phase up to temperature $T=\infty$ in 2D (critical $T_{c}$, i.e. finite-$T$ phase transition, in 3D) for any non-zero density of $\mathbb{Z}_{2}$ charges. Further, using quantum Monte Carlo we demonstrate that the POPs reproduce the square lattice Fradkin-Shenker phase diagram at $T=0$ and explore the phase diagram at $T>0$. The correlation length exponent coincides with the one of the 3D Ising universality class and we determine the POP critical exponent characterizing percolation. Our proposed POPs provide a geometric perspective of confinement and are directly accessible to snapshots obtained in quantum simulators, making them suitable as a probe for quantum spin liquids.
Introduction.—Lattice gauge theories (LGTs) have been widely studied in the fields of high-energy Wilson (1974), condensed matter Wegner (1971); Fradkin and Susskind (1978a); Kogut (1979) and biophysics Lammert et al. (1993). In 1979, Fradkin and Shenker proved in their groundbreaking work Fradkin and Shenker (1979) the existence of two phases in their model, where $\mathbb{Z}_{2}$ charged particles are confined or deconfined, respectively. Ever since, researchers have found intimate connections of $\mathbb{Z}_{2}$ LGTs to other physical effects, including topological order Wen (2007); Verresen et al. (2022), quantum spin liquids Read and Sachdev (1991); Sachdev (2019) and even quantum information Kitaev (2003). In the light of quantum simulation, LGTs with finite-dimensional local Hilbert spaces, e.g. $\mathbb{Z}_{2}$ LGTs and quantum link models Wiese (2013), gain particular attention because of their experimental feasibility Halimeh et al. (2023).
@@
Site percolation theory was first introduced in works by Flory and Stockmayer in the 1940s Flory (1941); Stockmayer (1944), where they studied the gel-point and the cross-linking of polymers. Nearly two decades later, bond percolation theory was introduced by Broadbent and Hammersley Broadbent and Hammersley (1957). They studied the random flow of a fluid through a medium and introduced the so-called Bernoulli percolation. A prime example of the use of percolation theory in physics is the random cluster (RC) model, which was introduced by Fortuin and Kasteleyn Kasteleyn and Fortuin (1969); Fortuin and Kasteleyn (1972); Fortuin (1972a, b). The RC model was successfully used to study many physical systems, e.g. the Potts model Grimmett (2006); Baxter (1982). More recently, percolation theory was applied in the context of quantum monopole motion in spin ice Stern et al. (2021), quantum error correction Hastings et al. (2014) and gauge theories Gliozzi et al. (2005).
In this Letter, we show that percolation can also serve as a non-local order parameter to probe confinement. In the confined phase, analogous to quarks in a confining potential, two charges are being held together by $\mathbb{Z}_{2}$ electric strings and form a $\mathbb{Z}_{2}$-neutral meson Homeier et al. (2023). In the deconfined phase, in our picture matter particles move essentially independently through the system and the $\mathbb{Z}_{2}$ electric strings form a global cluster spanning over the entire lattice, see Fig. 1. Charges attached to this cluster are no longer constrained to $\mathbb{Z}_{2}$ neutral hadrons. We study a classical and a quantum model using classical (quantum) Monte Carlo (MC/QMC). We show that the phase boundary in the celebrated Fradkin-Shenker model (extended toric code) Fradkin and Shenker (1979) can be reproduced by percolation-inspired order parameters (POPs), paving the way for the application of POPs in related models.
@@
Classical $\mathbb{Z}_{2}$ LGT.— The complex nature of percolation in $\mathbb{Z}_{2}$ LGTs emerges from the local $\mathbb{Z}_{2}$ symmetry generator
@@
To probe this transition, we define two POPs. Percolation probability is the probability that a cluster percolates, i.e. winds around the system and connects to itself, see Sup Sec. I.1. In the literature, this concept is sometimes also referred to as the wrapping probability Feng et al. (2008) or the winding number Banerjee et al. (2022). Percolation strength extends the percolation probability to include information about the size of the percolating cluster. When the system is not percolating, the percolation strength is zero. When the system is percolating, the percolation strength is defined as the number of strings in the largest string cluster divided by the total number of links Homeier et al. (2023). We probe the transition on a periodic square lattice with system sizes $L^{2}=20^{2},30^{2},40^{2},50^{2}$. In the following we distinguish the cases for (i) zero matter density $d=0$ and (ii) non-zero matter density $d=N/L^{2}>0$.
@@
Quantum $\mathbb{Z}_{2}$ LGT: extended toric code.—We start with a generalization of Hamiltonian (Percolation as a confinement order parameter in $\mathbb{Z}_{2}$ lattice gauge theories) by adding magnetic fluctuations $\propto J$ and dynamical matter $\propto\lambda$:
@@
At $h=\lambda=0.2$ (deconfined, topological ground state), we find a percolating phase at low but finite temperatures and a non-percolating phase at higher temperatures. We trace this back to the bulk gap, where due to exponential suppression of thermal matter excitations the topological ground state stays robust against a non-zero temperature in finite-size systems, see Sup Sec. II.2. At temperatures well above the bulk gap, we observe a clear decrease in the percolation strength with higher system sizes. We conjecture that the percolation probability approaches zero in the TDL for any $T>0$. We were unable to provide a good estimate of the Fredenhagen-Marcu order parameter for finite temperatures due to high noise in the data. In the literature, confinement at finite temperature is also probed using Polyakov loops Banerjee et al. (2024) or by explicitly violating the gauge symmetry Banerjee et al. (2013).
@@
Discussion and outlook.—We have introduced new POPs to probe the confinement of matter excitations in $\mathbb{Z}_{2}$ LGTs. We have applied it to study a classical model where we demonstrated the substantial influence of matter and the lattice geometry on the existence of a percolating, i.e. deconfined, phase. Further, using snapshots generated from continuous-time QMC, we demonstrated that the phase boundaries of the square lattice Fradkin-Shenker model, one of the textbook examples featuring confinement and topological order, can be reproduced by our POPs. The correlation length critical exponent falls into the 3D Ising universality class. We also demonstrated the potential of the POPs at finite temperature. Another task, left for future research, will be to check for a first-order percolation transition across the self-duality line at the tip of the deconfined phase.
While we only discuss $\mathbb{Z}_{2}$ LGTs on the square and the cubic lattice, we also envision future applications of the POPs in other LGTs, where different interactions, lattice geometries or higher symmetries can be realized. Percolation was already used in the context of quantum error correcting codes Botzung et al. (2023). Since percolation can be directly measured from snapshots in the electric field basis, the order parameter is readily accessible to state-of-the-art quantum simulators Zohar et al. (2017); Barbiero et al. (2019); Schweizer et al. (2019); Homeier et al. (2021, 2023); Mildenberger et al. (2022); Halimeh et al. (2023), and extends the geometric perspective on confinement beyond one-dimensional systems Kebrič et al. (2021, 2023).
We thank M. Aidelsburger, D. Banerjee, G. Dünnweber, M. Kebrič, F. Palm, G. Semeghini, T. Wang and H. Zhao for fruitful discussions. This research was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement no 948141) — ERC Starting Grant SimUcQuam, and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868 and via Research Unit FOR 2414 under project number 277974659. L.H. acknowledges support from the Studienstiftung des deutschen Volkes.
@@ I.2 Grand Canonical Monte Carlo Sampling
^$\hat{\mathcal{H}}_{\mathrm{gc}}=-h\sum_{l}\hat{\tau}^{x}_{l}-\mu\,\sum_{\bm{j}}\,\frac{1}{2}\,\biggl{(}1-\prod_{l\in+_{\bm{j}}}\hat{\tau}^{x}_{l}\biggr{)}.$ (S5)
@@ II.2 Extended toric code QMC
$h$-sweeps.— We set $T=1/L$, $\lambda=0,0.1,0.2,0.3$ and $L^{2}=20^{2},25^{2},30^{2}$. To extract the critical electric field $h_{c}$ and the critical exponents $\nu$ (correlation length) and $\beta$ (percolation strength $P$) we first perform a crossing point analysis of the percolation strength Binder cumulant
@@
$\lambda$-sweep.— We set $T=1/L$, $h=0.2$ and $L^{2}=10^{2},20^{2},30^{2}$. We find Binder cumulant crossing points around $\lambda\approx 0.4$ that shift toward lower $\lambda$ with increasing system size, see Fig. S5c. The gap consists of the bulk gap $\Delta_{\mathrm{B}}$ and the finite-size gap $\Delta_{L}$. Because of the self-duality, we know that the true phase transition where the gap closes is at $\lambda_{c}\sim 0.33$. However, in finite-size systems (i) the gap never completely closes and (ii) the finite-size ground state can appear topological because of the finite-size gap. We conjecture that the true critical value of $\lambda_{c}\sim 0.33$ is reached in the TDL and the strong finite-size effects are caused by a large finite-size gap. In Fig. S5b, we show the Fredenhagen-Marcu order parameter
@@
Figure S6: Ising mapping of the pure gauge model. We illustrate the mapping of the pure gauge model Hamiltonian (S11) to the TFIM on the dual lattice. In panel (a) we show two configurations of electric strings and the corresponding spin configuration on the dual lattice. The electric strings form closed loops. The electric field on a link $l$ depends on the dual lattice link between $\bm{i}^{\prime}$ and $\bm{j}^{\prime}$ that intersects $l$. By flipping a spin on a dual lattice site $\bm{j}^{\prime}$, all electric fields in the surrounding plaquette $\square_{\bm{j}^{\prime}}$ are flipped. A percolating string corresponds to a domain wall of spins since closed string loops necessarily have spins of the same orientation along their perimeter. Due to Gauss’s law, the star term $\propto\mu$ is constant, see panel (b). Adapted from Wegner (2014).
$T$-sweep.— We set $h=\lambda=0.2$ and $L^{2}=10^{2},20^{2},30^{2}$. We find a crossing-point in the percolation strength Binder cumulant at $T>0$ and a percolating phase extending into $T>0$, see Fig. S5d. However, the topological order is known to break down for $T>0$. We trace this behavior back to the bulk gap $\Delta_{\mathrm{B}}$. For $T<\Delta_{\mathrm{B}}$ the density of thermal excitations is non-zero in the TDL, however, it is exponentially suppressed and thus not visible in finite-size systems. After all, this is the reason why we can gain insights into the ground state using finite-$T$ QMC schemes. While the true ground state for $T>0$ is not topological, it appears like the topological ground state at $T=0$ in finite-size systems. This reasoning is supported by the estimated order-of-magnitude of the bulk gap $\Delta_{\mathrm{B}}\simeq J$ at $h=\lambda=0.2$ which matches the temperature scale below which we observe the percolating (deconfined) ground state features. We were unable to provide a good estimate of the Fredenhagen-Marcu order parameter for finite temperatures due to high noise in the data.
@@ III Mapping the pure gauge theory to the 2D transverse-field Ising model
Since we have no matter, the $\mathbb{Z}_{2}$ electric strings can only form closed loops. If we define Ising spins $\hat{s}^{z}\in\{\pm 1\}$ on the dual lattice sites, a string loop on an elementary plaquette can be thought of as one spin with $\hat{s}^{z}=+1$ in a background of spins with $\hat{s}^{z}=-1$. The operator $\hat{s}^{x}$ flips the spin $\hat{s}^{z}$ in the dual lattice and thus the electric fields on the surrounding plaquette in the LGT: it creates/destroys string loops. This intuition can be formalized with the definition of the dual-lattice site variables: